At a temperature of 27∘C. two identical organ pipes produce notes of frequency 140 Hz. If the temperature of one pipe is raised to 57.75∘C. the number of beats produced per second is
Ans: 7 beats/s
1
Expert's answer
2020-03-16T13:15:23-0400
"v=\sqrt{\frac{\gamma RT}{M}}"

"\frac{v_2}{v_1}=\sqrt{\frac{T_2}{T_1}}"

"\frac{f+x}{f}=\sqrt{\frac{T_2}{T_1}}"

"\frac{140+x}{140}=\sqrt{\frac{273.15+57.75}{273.15+27}}"

"x=7\frac{beats}{s}"